Suzuki had a total of 262 hits. He hit three triples than home runs, and he hit three times as many doubles as home runs. Suzuki also hit 45 times as many singles as triples. Set up 4 equations using the information given. s=single, d=double, t=triple, h=home run
Answer by macston(5194)

You give insufficient information to solve the problem. Following is a full version of the problem, but in the future, please provide the full problem. You will get a more timely response. I just happened to have a copy of this one:
(If this is not your problem, please disregard and re-post the full problem)
.
. Ichiro Suzuki holds the American League record for the most hits in a single baseball season. In 2004, Suzuki had a total of 262 hits for the Seattle Mariners. He hit three fewer triples than home runs, and he hit three times as many doubles as home runs. Suzuki also hit 45 times as many singles as triples. Find the number of singles, doubles, triples and home runs hit by Suzuki during the season.
.
You did not indicate how the problem was to be solved, so simplest method:
S=singles=45T=45(H-3)
D=doubles=3H
T=triples=H-3
H=home runs
S+D+T+H=262 Substitute for S,D, and T in terms of H
45(H-3)+3H+(H-3)+H=262 Solve for H.
45H-135+3H+H-3+H=262
50H-138=262
50H=400
H=8 . . ANSWER 1: Suzuki hit 8 home runs.
T=H-3=8-3=5 ANSWER 2: Suzuki hit 5 triples.
D=3H=3(8)=24 ANSWER 3: Suzuki hit 24 doubles.
S=45T=45(5)=225 ANSWER 4: Suzuki hit 225 singles.
CHECK:
S+D+T+H=262
225+24+5+8=262
262=262